Determine whether or not the following points are collinear.
(-2,-1),(5,5), and (7,9)

Answer by stanbon(75887)

Determine whether or not the following points are collinear.
(-2,-1),(5,5), and (7,9)
They are if pairwise they have the same slope:
slope = (5--1)/(5--2) = 6/7
slope = (9--1)/(7--2) = 10/9
slope = (9-5)/(7-5) = 4/2=2
----------------
I guess not.
Cheers,
Stan H.